• The short run production Q = L^3-2L^2-4L+40 determine the number labourers should be used to maximize production.
1
Expert's answer
2021-07-30T12:36:01-0400

"Q = L^3 \u2212 2L^2 \u2212 4L + 40"

To determine the number of laborers that would maximize production, differentiate production function wrt to L

"\frac{∂Q}{∂L} = 3L^2 − 4L − 4"

Set this equal to

"\frac{∂Q}{∂L} = 3L^2 − 4L − 4=0"


"3L^2 + 2L - 6L - 4 = 0\\

(3L + 2)L - 2(3L + 2) = 0\\

(L − 2)\times(3L + 2) = 0"

"L^* = 2 \space and\space L^* = \frac{-2}{3}"


Since Labor cannot be negative, Hence L* = 2

Hence, L* = 2 maximises production.